A hacker is trying to guess someone's password. The hacker knows (somehow) that the password is 4 digits long, and that each digit could be a number between 0 and 4. Assume that the hacker makes random guesses. What is the probability that the hacker guesses the password on her first try? Round to six decimal places.

Probability = _______

** Enter an integral or decimal number, accurate to at least 4 decimal places.
Answer #1

total number of passwords possible = 54 = 625

P(hacker guesses the password on her first try) = 1 / 625 = 0.0016 (ans)